If the hypotenuse of a right triangle is 30 cm long and one leg is twice as long as the other, how long are the legs of the triangle?
I think I'm suppose to set it up 30^2=a^2 +2b^2, but I'm not sure. And if it is, I'm not sure how to solve for both a and b. Thanks

Answer by mananth(16949)

if one leg is x
the other leg will be 2x
...
x^2+(2x)^2=30^2
x^2+4x^2=900
5x^2=900
x^2=900/5
x^2=180
x= sqrt 180 = 13.4 cm
other leg will be 2*13.4 = 26.8 cm